OpenStudy (anonymous):

At which point does the line y = 2 - x intersect the circle (x - 3)2 + y2 = 25?

OpenStudy (anonymous):

first write both the graphs in form P(x)=0

OpenStudy (anonymous):

sorry, solve both of them for y

OpenStudy (anonymous):

y=2-x \[y^2=25-(x-3)^2\] \[y=\pm \sqrt{25-(x-3)^2}\]

OpenStudy (anonymous):

\[2-x=\pm \sqrt{25-(x-3)^2}\]

OpenStudy (anonymous):

write this equation as two: \[2-x=\sqrt{25-(x-3)^2}\] and \[2-x=-\sqrt{25-(x-3)^2}\]

OpenStudy (anonymous):

solve each of them as usual

OpenStudy (anonymous):

\[(2-x)^2=25-(x-3)^2\] \[4-4x+x^2=25-x^2+6x-9\]

OpenStudy (anonymous):

\[2x^2-10x-12=0\] \[x^2-5x-6=0\] \[(x-6)(x+1)=0\] \[x={6, -1}\]

OpenStudy (anonymous):

do the same hing with second equation and check for extraneous solutions

OpenStudy (anonymous):

I got 3,4